OpenStudy (anonymous):

LAST QUESTION!!: Using complete sentences, explain the two ways to simplify sqrt180 and provide the simplified form.

OpenStudy (anonymous):

y once more u asking this?????????

OpenStudy (anonymous):

do it by urself

OpenStudy (anonymous):

Don't be rude sheg. This is a place to ask questions. Tyndale, I'm not sure what the question is asking. The square root can be written like so: \[\sqrt{x}=x^{1/2}\] so \[\sqrt{180}=180^{1/2}\] But that's not really a simplification. I think what the question is asking is something along the lines of http://answers.yahoo.com/question/index?qid=20110614002907AA9e3Pr